how to find the average of 3 mixed numbers (fractions)
1
Expert's answer
2015-05-20T10:24:11-0400
You should sum them up and then obtained result devide by three. Obtained then number is an arithmetic average of initial three numbers.
So the main feature here is to sum up mixed numbers. Actually its not a problem at all. For example, Let's add 2& 1/3 and 5& 4/7:
2 1/3+ 5 4/7=2+5+1/3+4/7=7+(1*7+4*3)/(3*7)=7+19/21=7 19/21